OpenStudy (anonymous):

HELP PLEASE!! Exercise 1-4 Calculate net income and stockholders' equity [LO2] [The following information applies to the questions displayed below.] Eagle Corp. operates Magnetic Resonance Imaging (MRI) clinics throughout the Northeast. At the end of the current period, the company reports the following amounts: Assets = $43,500; Liabilities = $20,500; Dividends = $2,140; Revenues = $13,100; Expenses = $8,200. 1) Calculate net income. 2) Calculate stockholders' equity at the end of the period.

OpenStudy (anonymous):

Net income is simply Revenues - Expenses. In this case 13100-8200=4900 For the second question, first we need to find out what the Shareholders Equity portion of balance sheet is. Assets= Liabilities+Shareholders Equity. From this we can determine S/E: 43500=20500+S/E S/E=23000 Now we need to consider the year's Net Income. From previous question we know that its 4900, however we have to take into account that dividends are payed out of this amount. Therefore the retained earnings after dividends are 4900-2140=2760. The total stockholder's equity is 23000+2760=25760. Let me know if you need any further clarification.